All of the following companies manufacture toasters. The following costs were incurred by each company during September:

Warner Co. Painted Trails Co. Catalina Inc. Sunrise Co.
Direct labor $12,000 $9,150 $72,200 $35,600
Rent 67,400 3,650 96,800 54,400
Utilities 2,700 1,400 16,300 7,900
Heating elements 86,900 42,000 100,900 64,800
Depreciation 5,900 9,700 28,000 37,800


Required:
1.
What is the total conversion cost for Warner Co.?
2. What is the total manufacturing overhead for Painted Trails Co.?
3. What is the total conversion cost for Catalina Inc.?
4. What is the total manufacturing overhead for Sunrise Co.?
Answer #1

1. Conversion cost = Direct Labor + Manufacturing Overhead

= 12000 + 67400 + 2700 + 5900

= 88000

2. Total manufacturing overhead = Rent + utilities + Depreciation

= 3650+1400+9700

= 14750

3. Conversion cost (check 1)

= 72,200+96,800+166,600 00 +28,000

= 213,300

4. Manufacturing overhead

= 54400+7900+37800

= 100100